The rate of effusion of Argon was measured to be 0.218 mol/s at a certain 

temperature. Calculate the rate of effusion for Helium. 


1
Expert's answer
2021-01-18T01:56:54-0500


We know that the rate of effusion and the molar mass of the gas related as


r∝(1/molar mass,M)

So,. r2/r1 = M1/M2

Or, r2 = r1×(M1/M2)____(1)

The rate of effusion of Argon,

r1 = 0.218 mol/s

Molar mass of argon, M1 = 49 g/mol



Let the rate of effusion of helium,

r2 = x mol/s

Molar mass of helium, M2 = 4 g/mol


Putting in the equation (1) we get


r2 = 0.218×(49/4)

= 2.6705 mol/s

Hence, rate of effusion of helium = 2.6705 mol/s